OpenStudy (anonymous):

write a recursive definition for the sequence 8, 6, 4, 2

OpenStudy (anonymous):

what would you respond if a 7th grader is trying to help you?

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

wait isn't a recursive definition basically a order for the numbers? similar to equations for graphing?

OpenStudy (anonymous):

no i am just curious....

OpenStudy (anonymous):

cz it looks like you are in high school or above?

OpenStudy (anonymous):

and i'm a freaking 7th grader

OpenStudy (goformit100):

t = a - (n-1)*2 where, t = term a = first term of the sequence n = place of term in the sequence

OpenStudy (goformit100):

@PrincessTianaB did you understand it ?

OpenStudy (anonymous):

aren't you just trying to find like a formula for the numbers? like so that you can continue the number chain or something?

OpenStudy (goformit100):

Yes @Molly_Qian

OpenStudy (anonymous):

not so much... how do you know which is term?

OpenStudy (goformit100):

Formula is: t = a - (n-1)*2 now take somew example

OpenStudy (goformit100):

*some

OpenStudy (anonymous):

\[t _{n}=t_{n-1} +b \] something like this?

OpenStudy (anonymous):

@goformit100

OpenStudy (p0sitr0n):

note that in this recursion, n starts at 1 (but could start at 0). up to you

OpenStudy (anonymous):

interesting....

OpenStudy (goformit100):

Hence you got the sequence you required: 8, 6, 4, 2 ....

OpenStudy (anonymous):

do you have to show work like that? it made it look harder than it is

OpenStudy (anonymous):

a1 = 8; an= an-1 -2 a1 = 8; an=an -1 + 2 a1 = 2; an= an-1 + 8 a1 = 2; an=an-1 + 6 are my options, so would it be the second option?

OpenStudy (goformit100):

to find first term of the sequence given by you, t = a - (n-1)*2 t_1 = 8 - (1 - 1)*2 = 8 to find first term of the sequence given by you, t = a - (n-1)*2 t_2 = 8 - (2 - 1)*2 = 6 to find first term of the sequence given by you, t = a - (n-1)*2 t_3 = 8 - (3 - 1)*2 = 4 to find first term of the sequence given by you, t = a - (n-1)*2 t_4 = 8 - (4 - 1)*2 = 2 here n is showing the position of term you want yo find.

OpenStudy (anonymous):

Would it be the second option since 8 is the first term and it adds two?

OpenStudy (anonymous):

i typed nothing \(\color{white}{I~think~it's~D}\)

OpenStudy (anonymous):

it adds two?

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

or minus two?

OpenStudy (goformit100):

Let me check

OpenStudy (anonymous):

stop it greta!

OpenStudy (anonymous):

Stop curing me out greta

OpenStudy (anonymous):

or \(\times -1 +6\)?

OpenStudy (anonymous):

what? @ees12

OpenStudy (anonymous):

^^^

OpenStudy (anonymous):

some guy named greta456 is sending me cursing messages

OpenStudy (anonymous):

\(\color{white}{I~still~think~it~is~D}\)

OpenStudy (goformit100):

@PrincessTianaB the options you gave is not clear. can you plz provide the screen short of the answers

OpenStudy (goformit100):

I mean options

OpenStudy (anonymous):

ohh okay. haha i was confused.

OpenStudy (anonymous):

@greta456 ? he is not online rn tho

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

What the heck, he's sending me messages right now.

OpenStudy (anonymous):

tag him. right on here or on os feedback or something. report him.

OpenStudy (anonymous):

\(\color{white}{I'm~pretty~sure~it's~D}\)

OpenStudy (anonymous):

he sent me this, A box contains 15 cards. There are 5 red cards labeled 1, 2, 3, 4, and 5. There are 5 blue cards labeled 1, 2, 3, 4, and 5. There are 5 green cards labeled 1, 2, 3, 4, and 5. One card is drawn from the box. What is the probability that it is a green card with an even number or a blue card?

OpenStudy (anonymous):

oops not that, I forgot to copy it, he sent me this https://www.youtube.com/watch?v=-yE_v6vN9UE

OpenStudy (anonymous):

ew.

OpenStudy (rational):

8, 6, 4, 2 you can get next term by subtracting 2 from present term, so the recursive definition can be \[a_{n+1} = a_n-2\\a_1 = 8\]

OpenStudy (anonymous):

@goformit100

OpenStudy (anonymous):

\(\color{white}{LAST~CHOICE}\)

OpenStudy (anonymous):

So it would be the first option, @rational ?

OpenStudy (anonymous):

https://www.youtube.com/watch?v=zqPN4UbnlWg

OpenStudy (anonymous):

He sent me that too.

OpenStudy (rational):

Somehow I am not able to see options ?

OpenStudy (rational):

Yes first option :)

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @ees12 some guy named greta456 is sending me cursing messages \(\color{blue}{\text{End of Quote}}\) please don't spamm this post where other users are trying to work you have " report" oor "block" option thanks

OpenStudy (anonymous):

Np barbie

OpenStudy (anonymous):

Thank you @rational (:

OpenStudy (goformit100):

yes they are correct